I'm having problems solving this algebra word problem.
A water tank has an inlet pipe and a drain
pipe. A full tank can be emptied in 30 minutes if the drain
is opened and an empty tank can be filled in 45 minutes
with the inlet pipe opened. If both pipes are accidentally
opened when the tank is full, then how long will it take to
empty the tank?

Answer by stanbon(75887)

A water tank has an inlet pipe and a drain
pipe.
A full tank can be emptied in 30 minutes if the drain
is opened and an empty tank can be filled in 45 minutes
with the inlet pipe opened.
If both pipes are accidentally opened when the tank is full,
then how long will it take to empty the tank?
------
Emptying pipe:
time = 1/2 hr/job ; rate = 2 jobs/hr
------
Filling pipe:
time = 3/4 hrs/job ; rate = 4/3 job/hr
-------------------------------------------
Together DATA:
time = x hr/job rate = 1/x job/hr
=====================================
Equation:
empty rate - fill rate = together rate
2 - 4/3 = 1/x
---
2/3 = 1/x
x = 3/2 hrs (time to fill the tank)
==========================
Cheers,
stan H.